Janice bought a 13 foot inflatable snowman to display in her yard during the holidays. She bought a 22 foot rope to anchor to the top of the snowman. How far away from the base of the snowman should she stake the other end of the rope so the snowman doesn't lean?
Answer by nerdybill(7384)

Apply Pythagorean theorem:
.
Let x = distance of stake
.
x^2 + 13^2 = 22^2
x^2 + 169 = 484
x^2 = 315
x = 17.75 feet